### v1

# Survey on Question Answering over Visually Rich Documents: Methods, Challenges, and Trends

Abstract
Using Large Language Models (LLMs) for Visually-rich Document Understanding (VrDU) has significantly improved performance on tasks requiring both comprehension and generation, such as question answering, albeit introducing new challenges. This survey explains how VrDU models enhanced by LLMs function, covering methods for integrating VrD features into LLMs and highlighting key challenges.
1 Introduction
Visually-rich documents (VRDs) combine complex information, blending text with visual elements like graphics, diagrams, and tables to convey detailed content effectively Ding et al. (2024a). Unlike traditional text documents, VRDs have two main features: text associated with typographic details (e.g., font, size, style, color), layout that organize information spatially, and visual elements, such as charts and figures, which enhance comprehension Huang et al. (2024a). These documents can be either native digital files (e.g., PDFs) containing searchable text and layout metadata, or scanned images requiring OCR to extract text and layout. Visually-rich Document Understanding (VrDU) is a rapidly evolving field at the intersection of computer vision and natural language processing, tackling both perception (document parsing, i.e. identification and extraction of objects within the document) and interpretation (analytical tasks on the document features, such as answering questions) Zhang et al. (2024c).
Current question-answering tasks on VrDU are not purely extractive; they require reasoning skills such as arithmetic, logical, and spatial reasoning, as well as counting, sorting, comparison, and multi-hop reasoning, which may involve combining multiple elements to answer a question, potentially spanning across modalities, pages, or even documents, as depicted in Table 1, illustrated by Figure 1.
Current approaches in document understanding have moved away from the traditional two-step pipeline of document parsing (converting the document into structured textual format) followed by interpretation (passing this structured text to a language model for answering questions) Huang et al. (2024b). Instead, they leverage a unified representation of the document that is directly processed by a LLM to tackle complex tasks. In this paradigm, the document parsing Zhang et al. (2024c) part becomes embedded within the model’s weights, with VrD encoders pretrained on document parsing tasks to learn representations that effectively capture all relevant document attributes Al-Shetairy et al. (2024), including contents (textual and visual), reading order, and specific layout information. We focus on this line of research in this paper.
To our knowledge, recent VrDU surveys focus on specific data types like charts Al-Shetairy et al. (2024) or tables and figures Huang et al. (2024b). The only surveys on whole documents (such as PDFs) are a catalog of models Ding et al. (2024b); Abdallah et al. (2024). More generally, no survey has thoroughly compared the different approaches to representing VrDs, specifically vision-only versus multi-modalities representations. So we provide a comprehensive analysis of how VrDU models interpret VrDs that often include multiple elements such as charts, tables, figures, and text, and span multiple pages (see Table 1), to answer questions. We break down the VrDU process into three key components: (1) Document representation – exploring encoding methods that incorporate text, layout, and visual data (Section 2) or vision-only approaches (Section 3), with a focus on multi-page (MP) representation (Section 4); (2) Embedding integration with LLMs – examining where and how document embeddings are incorporated (Section 5); and (3) Efficient training of these VrDU models, focusing on the pretraining approaches for modalities alignment (Section 6).
Document Characteristics (per document) Questions Characteristics
Datasets type #Pages #Tokens #Tab #Fig Crosspage Unans. Crossdoc #Regions Answer length
VisualMRC 2021 Wikipedia pages 1.0 151.46 ? ? 9.55
DocVQA 2021a Industry Documents 1.0 182.8 ? ? 2.43
InfographicVQA 2021b Posters (Canva, …) 1.2 217.9 ? ? 1.6
TAT-DQA 2022 Annual Companies Reports 1.3 550.3 >1 ? 3.44
MP-DocVQA 2023b Industry Documents 8.3 2026.6 ? ? 2.2
DUDE 2023 archives, wikimedia, … 5.7 1831.5 ? ? ✓(2.1%) ✓(12.7%) 3.4
SlideVQA 2023 Slides from Slideshare 20.0 2030.5 ? ? ✓(13.9%) $\approx$1
MMLongBenchDoc 2024c ArXiv, Reports, Tutorials, … 47.5 21214.1 25.4% 20.7% ✓(33.0%) ✓(22.5%) 2.8
M3DocVQA 2024 Wikipedia pages 12.2 ? ? ? ? ✓(2.4k) ?
M-LongDoc 2024 ArXiv, Manuals, Reports 210.8 120988 71.8 161.1 180.3
MMDocBench 2024 Multi 1.0 ? ? ? 2.61 4.1
Table 1: Overview of Question-Answering VrDU datasets on PDFs or PowerPoint documents, summarizing Document Characteristics (e.g., average pages, tokens, tabs, figures per document) and Question Characteristics (e.g., presence of questions requiring cross-pages or cross-documents information, unanswerable questions, and average answer length). #Region refers to the number of regions identified for answering questions in datasets with coordinate annotations. Underlined datasets are standard benchmarks used for model comparison in Table 5.
2 Encoding VrDs from structured information
VrDs can be represented through three distinct but interconnected modalities: text and layout, derived from native digital formats or OCR extraction, and the overall visual appearance of the document, obtained by generating a screenshot of the document page. The most important layout features are bounding boxes around text and structural elements (e.g., tables). The visual modality captures the document page appearance, encompassing its overall structure and visual context of the document as a whole. The main problematic in VRD encoding is to represent and merge the information coming from these three distinct modalities. Table 2 summarizes encoder models VrD and that we describe in this section.
Model $\mathbf{E_{Text}}$ $\mathbf{E_{Vis}}$ $\mathbf{E_{Pos}}$ $\mathbf{E_{Cross}}$ $\mathbf{D_{Text}}$ MP
Interaction of text and visual features within self-attention after modalities concatenation
LayoutLMv2 2022 UniLMv2 ResNeXt-101-FPN emb. tables + attn bias transformer
LayoutXLM 2021 XLM-R ResNeXt-101-FPN emb. tables + attn bias transformer
UNITER 2020 BERT Faster R-CNN emb. tables (7D) transformer
LayoutLMv3 2022 RoBERTa ViT attn bias transformer
DocFormerv2 2023 T5 encoder ViT emb. tables. T5
GRAM 2024 DocFormerv2(2023) DocFormerv2(2023) emb. tables DocFormerv2(2023)
LayoutLLM 2024 LayoutLMv3(2022) LayoutLMv3(2022) LayoutLMv3(2022) Llama-7B
DocLayLLM 2024 LayoutLMv3(2022) LayoutLMv3(2022) LayoutLMv3(2022) Llama3-8BInstruct
Interaction of text and visual features within cross-attention
DocFormer 2021 LayoutLM(2019) ResNet50 emb. tables visual-spatial attn transformer
SelfDoc 2021b Sentence BERT Faster R-CNN emb. tables intra&inter-modal attn transformer
ERNIE-Layout 2022 BERT Faster R-CNN emb. tables Disentangled attn (2021b) transformer
HiVT5 2022b T5 encoder DiT (2022) emb. tables VT5 encoder VT5 decoder
DocTr 2023 LayoutLM(2019) DETR (2020) special tokens Deformable DETR (2021) LayoutLM
InstructDr 2024 FlanT5 encoder CLIP VIT-L/14 emb. tables Document-Former FlanT5
RM-T5 2023 T5 encoder DiT (2022) emb. tables RMT (2022) T5 decoder
Arctic-TILT 2024 T5 encoder U-Net (per RoI) attn bias Tensor Product T5
Summing aligned text and visual features via ROI-pooling
TILT 2021 T5 encoder U-Net attn bias T5
Pramanik et al. (2022) Longformer ResNet50 + FPN sinusoidal emb. transformer
UDOP 2023 T5 encoder MAE encoder attn bias T5&MAE decoder
Interaction of text and visual features with hierarchical / graph architectures
PICK 2020 transformer CNN GNN BiLSTM + CRF
GraphDoc 2022 SentenceBERT SwinT + FPN emb. tables GAT (2018) transformer
LAMPreT 2021 BERT CNN Hierarchical fusion transformer
Table 2: Comparison of VrDU models handling the three modalities (T+L+V), detailing encoding of text $\mathbf{E_{Text}}$, visuals $\mathbf{E_{Vis}}$, and position $\mathbf{E_{Pos}}$, fusion layers $\mathbf{E_{Cross}}$, decoder $\mathbf{D_{Text}}$, and multi-page (MP) support ✓.
2.1 Integrating the Layout information
The positions and sizes of elements within a document can vary in granularity, from individual tokens Łukasz Garncarek et al. (2021) to larger blocks like cells, tables, images, or paragraphs Li et al. (2021a, b). This layout information can be represented within VrDU models in three ways: through positional embeddings, attention bias, or directly within the text, as special tokens.
The simplest approach, which does not require any architectural change, is to include layout information as special tokens, directly within the text. Lyrics Li et al. (2023c) and ViTLP Mao et al. (2024) add special tokens to represent layout information alongside text tokens. In Lyrics Li et al. (2023c), each word is tagged with its spatial coordinates using the $<Bbox>$ token, indicating its position within the document. ViTLP Mao et al. (2024) uses a $[LOC]$ token for the same purpose. The global text-layout sequence is based on an extended vocabulary $\hat{V}=V\cup[\text{LOC}]$, where $V$ is the original text vocabulary. This approach not only increases the sequence length, overloading the model’s context window, but also limits the ability to capture complex spatial interactions between elements in the document.
This is why the VrDU community has focused on developing optimal methods to incorporate spatial information of tokens within documents. One way is to extend the 1D positional encoding of tokens in transformers to 2D (see Table 2) by embedding the spatial coordinates $(x,y)$ of each token’s bounding box. For example, LayoutLM Xu et al. (2019) embeds the $x$ and $y$ coordinates separately and sums them. DocFormer Appalaraju et al. (2021) further includes embeddings for the bounding box dimensions (height and width), while UNITER Chen et al. (2020) adds an embedding for the area of each bounding box. These embeddings can be learned or fixed (function-based, e.g., sinusoidal Hong et al. (2021)). Given the observation that 2D dependencies might vary across layers, some methods make the layout embeddings layer-specific Appalaraju et al. (2021). However, these approaches ignore the relative position of tokens within the document.
To address this, some works extend the sequential positional attention bias used in seq-to-seq transformers, added before the softmax, to a 2D spatial context. This 2D positional bias represents the horizontal and vertical distances between tokens within the document Xu et al. (2022); Powalski et al. (2021). Fixed distance biases in attention are limited, as token semantics, like "total" in tables, often dictate specific spatial interactions beyond mere positional proximity. To ensure that the model pays particular attention to tokens located at the same horizontal position (as a token "total" for example), ERNIE-Layout Peng et al. (2022) introduces three relative position attention biases (disentangled attention), capturing respectively how the semantic meaning of a token interacts with its sequential, horizontal and vertical relative distance to the other token. FormNet Lee et al. (2022) goes further in this direction by allowing more complex interactions, using functions that combine semantic and position information, between the tokens semantics and relative positioning when computing the attention bias. This allows for instance to identify a label token positioned above or to the left of its corresponding value token.
For complex document layouts, methods that use positional biases in attention—especially semantic-based relative positional biases like the disentangled attention Peng et al. (2022) or the Rich Attention Lee et al. (2022)—seems to be more effective, at least on the Information Extraction tasks they have been compared on Jaume et al. (2019); Park et al. (2019).
2.2 Integrating the visual information
In all the reviewed works, the visual modality is a set of visual “tokens” (vectors), generated by a visual encoder. Initially based on CNNs Xu et al. (2022), these encoders have transitioned to Visual Transformers (ViTs) Huang et al. (2022).
Fusing text and visual features for unified document encoding is challenging due to the differences between visual and text tokens (see Table 2).
Local modality alignment refers to aligning text and visual features specifically within localized regions of the document, focusing solely on the text and visual attributes from those regions. These regions can be either inferred using visual information, i.e. determined by an object detection module Carion et al. (2020); Ren et al. (2016) or determined by the textual information, i.e. considering the bounding boxes of text tokens Powalski et al. (2021). A simple method to locally align modalities involves summing the two representations per region Powalski et al. (2021). Note that regions without associated text only have a vision-only representation Tang et al. (2023).
Another line of works process differently visual and textual tokens representations. SelfDoc Li et al. (2021b) uses two cross-attentions for local alignment: from the visual to textual tokens and vice-versa, e.g. allowing the textual semantic representation to be contextualized by visual information such as color, bold elements, and position. For example, a large, bolded, centered text block is likely to serve as a title or header. By incorporating these visual cues, the model refines the semantic representation of text, ensuring that its meaning is informed by its visual context within the document. Rather than relying on costly cross-attention for modality fusion and interaction, Arctic-TILT Borchmann et al. (2024) introduces a lightweight attention mechanism after the transformer feed-forward layer to integrate visual information using a learnable role bias for text tokens, inspired by the TP-Attention Schlag et al. (2020).
Global modality alignment involves considering both the visual and textual features of the entire document rather than on specific regions. A simple method to align those modalities globally involves concatenating them Xu et al. (2022). A transformer encoder then allows interaction through standard self-attention mechanisms Appalaraju et al. (2023); Huang et al. (2022). On the document scale, visual and textual tokens representations can also be processed differently. DocFormer Appalaraju et al. (2021) uses cross-attention for textual queries to guide the integration of visual information, and vice versa.
Finally, visual and textual tokens can interact by viewing the document as a graph, built thanks to the bounding boxes of elements like paragraphs, tables or figures, each representing a node. Graph-Based Relationship Modeling employs graph neural networks (GNNs) to capture relationships between modalities in documents Yu et al. (2020). Various graph-based approaches differ in how they aggregate neighbor information between nodes: Techniques include convolution (Graph Convolutional Networks) Yu et al. (2020), attention (Graph Attention Networks) Li et al. (2023d), and other aggregation methods Lee et al. (2022). While graph-based modeling is particularly effective at capturing long-range dependencies and contextual relationships in documents, it remains limited in facilitating interactions between the semantic of modalities, since modalities interaction is only layout-based.
3 Vision-Only Encoding of VrDs
Model Res. $\mathbf{E_{Vis}}$ $\mathbf{P_{E_{V}\to D_{T}}}$ $\mathbf{D_{Text}}$ MP
Encoder: HR image – Decoder: Tiny Decoder
DONUT Kim et al. (2021) 2560x1920 SwinT (2021a) MLP BART
DESSURT Davis et al. (2022) 1152x768 Attn-Based CNN MLP BART with Swin attn
Pix2Struct Lee et al. (2023a) 1024x1024 ViT MLP BART
SeRum Lee et al. (2023b) 1280x960 SwinT 2021a MLP mBART
Kosmos2.5 Zheng et al. (2023) 224x224 Pix2Struct 2023a’s ViT Perceiver Resampler Transformer
Encoder: LR image – Decoder: LLM
LLaVAR Zhu et al. (2023) 336x336 CLIP VIT-L/14 MLP Vicuna13B
Unidoc Nan et al. (2023) 336x336 CLIP VIT-L/14 MLP Vicuna13B
mPLUG-DocOwl Ye et al. (2023b) 224x224 CLIP VIT-L/14 Visual Abstractor Llama-7b
QwenVL Bai et al. (2023a) 448x448 CLIP-VIT-G/14 Cross-attn layer Qwen-7b
Encoder: HR image – Decoder: LLM thanks to HR image in subimages division (Section 3.2)
SPHINX Chen et al. (2023c) 1344x896 VIT & ConvNext & MLP Llama2-7B
DINO & QFormer
UREADER Liu et al. (2023b) 2240x1792 CLIP ViT-L/14 MLP Vicuna13B
Monkey Yin et al. (2023) 1344x896 CLIP Vit-BigG Perceiver Resampler Qwen-7B
TextMonkey Yin et al. (2024) 1344x896 CLIP Vit-BigG Shared Perceiver Resampler Qwen-7B
mPLUG-DocOwl1.5 Li et al. (2024a) 2560x1920 EVA-CLIP H-Reducer Llama-7b + MAM
LLaVA-UHD Jiang et al. (2024) 672x1088 CLIP-ViT-L Shared perceiver Resampler Vicuna-13B
InternLMXC2-4KHD Hu et al. (2024b) 3840x1600 CLIP-ViT-L PLoRA matrix InternLM2-7B
Idefics2 Beurton-Aimar et al. (2024) 980x980 SigLIP-SO400M MLP Mistral-7B-v0.1
TextHawk Wang et al. (2024) 1344x1344 SigLIP-SO Perceiver Resampler InternLM-7B
TokenPacker Li et al. (2024c) 1344x1344 CLIP-ViT-L TockenPacker Vicuna-13B
mPLUG-DocOwl2 Zhang et al. (2024d) 504x504 EVA-CLIP H-Reducer+DocCompressor Llama-7b + MAM
Encoder: HR image – Decoder: LLM thanks to adaptation of ViT to capture fine-grained details (Section 3.1)
DocPedia Chen et al. (2023a) 2560x2560 SwinT 2021b MLP Vicuna-13B
LLaVA-PruMerge Liu et al. (2023a) 336x336 CLIP-ViT MLP Vicuna13B
CogAgent Chen et al. (2023b) 1120x1120 EVA2-CLIP & CogVLM Cross-attn layer & MLP Vicuna-13B
Vary Li et al. (2023a) 1024x1024 ViTDet & CLIP-ViT-L MLP Qwen-7B
Mini-Gemini Lu et al. (2024) 2048x2048 ConvNeXt & ViT-L/14 MLP Mistral-7B
LLaVA-HR Li et al. (2024b) 1024x1024 CLIP-ConvNeXt & ViT-L MLP & MR-Adapter Llama2-7B
TinyChart Xie et al. (2024a) 768x768 SigLIP MLP Phi-2
HRVDA not specified (2024) 1536x1536 SwinT (2021b) MLP Llama2-7B
DocKylin Zhang et al. (2024a) 1728x1728 SwinT (2021b) MLP Qwen-7B
Table 3: Comparison of vision-only VrDU models, detailing the input image resolution (Res), visual encoding $\mathbf{E_{Vis}}$, vision-to-text projection $\mathbf{P_{E_{V}\to D_{T}}}$, decoder $\mathbf{D_{Text}}$, and multi-page (MP) support ✓.
In the previous section, we discussed many different techniques to integrate visual and textual information. These models however remain complex pipelines because they need to deal with the different modalities.
Many recent works consider VrDs as images, which brings the advantage of dealing with only a single modality, relying on a (textual) decoder to handle different tasks. A summary of this type of model whose architecture we detail below is provided in Table 3.
Such approaches demand a highly capable visual encoder to capture all textual, layout, and visual details within the document. However, ViTs themselves are not capable to capture fine details within documents like text Anonymous (2024). Indeed, in ViTs, the visual input (e.g., a document page) is divided into fixed-size patches, each becoming a "vision token" (e.g., 14x14 or 16x16 pixels). If patches are too large, they may cover too much content, like multiple lines or text fragments, and miss fine details. Using smaller patches or increasing the image resolution creates more patches, enabling the model to capture finer details and better encode the document’s visual and textual content Lee et al. (2023a), but at the cost of efficiency.
Indeed, ViTs have a maximum context size (number of patches) they can manage Lee et al. (2023a). This is why research in vision-only VrDU focuses on architectural modifications to ViTs to enable the processing of high-resolution images (Section 3.1). A simpler alternative is to use a set of pre-trained ViTs, each handling a different part of the image, thereby allowing the processing of high-resolution images more efficiently (Section 3.2). In this case, the local of a global context might be harmful and needs to be tackled.
3.1 Architectural changes to ViT
A number of approaches leverage CNN architectures, which capture local information more efficiently than ViTs due to their intrinsic design based on convolutions, exploiting equivariant biases inherent to images. Dhouib et al. (2023) proposes a sequential architecture combining CNN and ViT components, where ConvNext blocks are used to extract local features, and their output is fed into a ViT for modeling global dependencies.
Due to the complexity of combining two networks, a CNN and a ViT, without losing information, other approaches Kim et al. (2021); Blecher et al. (2023) draw inspiration from the local window mechanism of CNNs and incorporate it into ViTs, enabling them to process numerous patches effectively. These approaches restrict attention to a local window of patches with a Swin Transformers Liu et al. (2021a), which applies self-attention within local windows, shifting these windows across layers to efficiently integrate cross-window information. However, Swin ViTs progressively reduce the resolution of the tokens through token merging steps, which decrease the number of tokens. DocPedia Feng et al. (2024) removes this downsampling step, enabling it to keep the full token resolution throughout the processing pipeline by leveraging the frequency domain rather than spatially merging patches as done in Swin. Indeed, decomposing an image into the frequency domain, such as using the Discrete Cosine Transform (DCT) Liu et al. (2022a), allows to process larger patches without loosing important high resolution information.
More recent approaches avoid introducing a local bias to ViTs, instead focusing on removing redundant information from ViT patches, as documents often contain a significant amount of redundancy, such as borders, whitespace or decorations. These methods either use attention scores from the self-attention mechanism to prune or merge tokens (e.g., Zhang et al. (2024b); Liu et al. (2023a); Zhao et al. (2023)) or employ unsupervised techniques like Dual-Center K-Means Clustering Zhang et al. (2024a) to compare tokens. TinyChart Zhang et al. (2024b) merges similar tokens at every ViT layer using methods like average pooling, while DocKylin Zhang et al. (2024a) employs similarity-weighted summation based on token cosine similarity ensuring that each token contributes proportionally to its relevance. HRVDA not specified (2024) uses probabilities from a content detector to guide merging, using a shallow segmentation network that focuses on pixel-level information, assigning probabilities to pixels and mapping them to corresponding patches to filter out low-relevance areas (e.g. whitespace or decorations) while preserving meaningful regions (e.g. text or tables). DocKylin Zhang et al. (2024a) goes further in the pixel-level direction of HRVDA not specified (2024), before the image tokenization, to identify redundant information: A preprocessing module (Adaptive Pixel Slimming) modifies the image itself, removing whitespace by detecting low-gradient areas (smooth regions lacking significant content). The identified redundant rows and columns are then cropped, resulting in a more compact and efficient document representation.
These approaches, which modify ViTs by adding local biases (e.g., via CNNs in hybrids or Swin ViTs) or using patch merging/selection, often result in information loss due to merging, pruning, or restricted representations. Recent works have explored pipelines with no modification of the ViT architecture, leveraging pretrained ViTs to process high-resolution images cut into slices, each ViT handling a specific portion of the image.
3.2 Several ViTs to process partitioned image
The idea of these methods is to slice a high-resolution image into subimages, each processed in parallel by an independent ViT. The resulting representations are combined for a unified document representation.
The way the original image is sliced into subimages is crucial to prevent information loss. Padding preserves the aspect ratio and prevents deformation Li et al. (2024c). Some approaches predict the optimal way to cut the original image, with pre-defined grid matching Ye et al. (2023a) and a score function predicting the best partition Xu et al. (2024), resulting in a varying amount of crop. Whatever the method, models need to maintain the continuity between the different subimages representations.
A first way to do so is through a transformer layer, which helps the model correlate local images to alleviate salient information loss due to cropping Ye et al. (2023a). Liu et al. (2024) introduces a Shifted Window Attention mechanism, enabling sliding window-based attention across subimage representations.
A more efficient approach to maintain continuity between subimages is to leverage a low-resolution document representation to guide the integration of subimages. Through a cross-attention layer, TokenPacker Li et al. (2024c), and later mPLUGDocOwl2 Zhang et al. (2024d), integrate the high-resolution representation of regions into the low-resolution representations using cross-attention, thus interpolating these low-resolution representations with its multi-level region cues treated as reference keys and values to inject their finer information to global image view.
To conclude on vision-only approaches, we think that slicing approaches using local information from cropped image regions to complement a low-resolution global view are promising, enabling compact and efficient representations with significantly fewer tokens while maintaining essential layout and semantic details Zhang et al. (2024d). We also think that techniques like adaptive pixel slimming Zhang et al. (2024a) are promising for VrDU, leveraging the inherent structure of documents, containing whitespace or decorations, to preserve all its content.
4 Encoding multi-pages documents
Multi-page documents are common in VrDU and are challenging for transformer-based architectures due to their length. Multi-page documents vary in lengths (e.g., 20 pages in SlideVQA Tanaka et al. (2023)), information density (e.g., 21214 in MMLongBench-Doc Ma et al. (2024c)), and cross-page informations (e.g. 2.1% in DUDE Landeghem et al. (2023)). To encode multi-page documents, recent approaches use retrieval-augmented generation (RAG) techniques Lewis et al. (2021) (Section 4.1). Other methods represent the document page by page (Section 4.2), while some rely on long-sequence processing techniques for page interactions (Section 4.3).
4.1 Retrieval Approach to multi-page
The retrieval approach to multi-page documents focuses on supplying to the VrDU decoder only the representation of pages with relevant information. Several levels can be used to identify the "oracle" element from the document: the retriever can either predict the entire oracle page Naidu et al. (2024); Faysse et al. (2024); Ma et al. (2024b); Cho et al. (2024) or focus on specific regions within the page, such as paragraphs or images that contain the relevant information for answering the question Xie et al. (2024b).
These approaches inherently limit either the interaction between pages or the interaction between modalities, which does not allow cross-page analysis Ma et al. (2024c). Other possibilities, described in the next sections, can be used for multi-modal encoders, but they are however the only viable option to enable vision-only models to handle multi-page inputs (see Table 5), since vision-only models encode a page with many tokens.
4.2 Query-biased approaches
HiVT5 Tito et al. (2023a), and later InstructDr Tanaka et al. (2024), encode each page of the document separately, with a specific learnable token added at the start of each page. HiVT5 Tito et al. (2023a) uses the specialized [PAGE] tokens to guide the encoder in summarizing each document page based on the given question, by processing separately each page with the question, encoding all the relevant information for the next processing step into the [PAGE] token. These [PAGE] tokens representations are then concatenated and passed to the decoder to generate the final answer. To our knowledge, the only vision-only model designed for multi-page input is mPLUG-DocOwl2 Zhang et al. (2024d), which compresses each page representation into 324 tokens and adds a page token for each page. In vision-only approaches, the token length of high-resolution images (i.e., document pages) is typically too large for LLMs to handle multi-page joint understanding, necessitating extreme compression of each page representation and thus degrading performance Zhang et al. (2024d).
However, query-biased approaches only allow limited cross-page reasoning, as the long sequence and diluted information across pages make it challenging to capture specific inter-page relationships Ma et al. (2024c), the page token being not leveraged effectively.
4.3 Efficient encoding of multi-pages
Inspired by the ETC Global-Local Attention mechanism Ainslie et al. (2020), GRAM Blau et al. (2024) enables global reasoning across multiple pages through a combination of page-dedicated layers, which apply self-attention within each page representation, and document-level layers, which focus exclusively on page token embeddings in their attention computations.
Another sparse attention approach is implemented by Arctic-Tilt Borchmann et al. (2024), employing a blockwise attention strategy limiting the attention to a chunk size, allowing to handle up to 500 pages (about 390k tokens, or 780 tokens per page on average). This method limits attention to a smaller, predefined neighborhood, reducing complexity from quadratic to linear while representing cross-page information.
An alternative to sparse attention for efficient multi-page documents processing is to use a recurrent network. RM-T5 Gupta et al. (2023) uses a Recurrent Memory Transformer (RMT) Gupta et al. (2022) to process multi-page documents sequentially, treating each page as part of a sequence. This allows the model to carry information across pages by utilizing hidden states from previous pages. The RMT selectively retains or forgets information, capturing essential details from each page for the next encoder, with all memory cells concatenated for the decoder to generate the final answer. However, the drawbacks of RNNs are inherited, such as the lack of parallelization and the limited possible interaction of two elements (here, pages) distant in the sequence.
Overall, we think that approaches that encode entire documents using sparse attention techniques, either global-local or blockwise, represent the future of the multi-page field, as they show great performance on cross-page reasoning Ma et al. (2024c) over retrieval ones.
5 Injecting VrD features into the LLM
Given their high quality generation abilities, large language models (LLMs) have been leveraged in VrDU (see Table 3). Integrating VrD features into an LLM decoder is not straightforward because it requires adapting the VRDs representation space to into an LLM-compatible representation without losing information, while preserving some computational efficiency.
5.1 Self-attention based approach
This self-attention approach Laurençon et al. (2024) consists in prepending the VrD representation to the prompt, allowing the model to process both VrD features with the prompt together in its self-attention layers. In such approaches, VrD features are projected to the LLM space via several approaches, and are optionally pooled into a shorter sequence.
Methods range from direct linear projection using a single layer to map VrD tokens to the expected input format of the language model Lee et al. (2023a), which projects tokens with minimal parameters, to convolutional ones Yang et al. (2023), allowing to reduce the dimension of the VrD representation, and to using learnable queries to retrieve the relevant VrD tokensLi et al. (2023b); Bai et al. (2023b).
Since interactions within visual tokens are already handled by the vision encoder in vision-only approaches, Ma et al. (2024a) modify the self-attention mechanism of the LLM by a Composite-Attention, removing interactions within the LLM within visual tokens; text tokens act as queries, with both visual and text tokens serving as keys and values.
These approaches are limited as they consider the raw tokens of the textual prompt at the same level, without distinguishing between their respective roles or significance.
5.2 Cross-attention based approach
In the cross-attention-based approach, VrD hidden states encoded by the VrDU encoder are used to condition the frozen LLM using freshly initialized cross-attention layers which are interleaved between the pretrained LLM layers Laurençon et al. (2024). Unlike self-attention, cross-attention approach enables a separate consideration of prompt and VrD tokens. Flamingo Alayrac et al. (2022) pioneered this approach with its Perceiver Resampler, which has since been adopted in various VrDU models (see Table 3).
An advantage of using cross-attention is that it allows to process longer sequences from the encoder, and thus to use only high-resolution representations. For instance, the CogAgent Hong et al. (2023) model employs a high-resolution encoder connected to the decoder through a cross-attention layer: The high-resolution hidden-states condition the LLM.
In other words, cross-attention approaches for integrating visual document (VrD) features into LLMs enable the query/prompt tokens to explicitly interact with VrD features, effectively leveraging the LLM’s capabilities. However, these methods require the introduction of many new parameters, as cross-attention layers are interleaved with the LLM’s architecture, significantly increasing the overall model size Laurençon et al. (2024). We describe this training process in the next section.
6 Training VrDs
VrDU require pre-training for modalities alignment in T+L+V approaches (Section 2), for adapting vision-encoders to text-centric tasks in V-only approaches (Section 3). When using LLMs, additional training is also needed for the interface between the encoder and the LLM.
Model Pretraining Tasks
DocFormer 2021 MVLM, LTR, TDI
TILT 2021 MVLM
LayoutXLM 2021 MVLM, TIA, TIM
SelfDoc 2021b MRP
Pramanik et al. (2022) MVLM, DTM, DSP
LayoutLMv2 2022 MVLM, TIA, TIM
LayoutLMv3 2022 MLM, MIM, WPA
GraphDoc 2022 MSM
ERNIE-Layout 2022 ROP, RRP, MVLM, TIA
UDOP 2023 MVLM, JTL, VTR, MIR
LAMPreT 2021 BOP, MBP, IFP
DocFormerv2 2023 TTL, TTG
GRAM 2024 None
HiVT5 2022b MVLM
DocTr 2023 MDM
Arctic-TILT 2024 MVLM
LayoutLLM 2024 MLM, MIM, WPA
Table 4: Pretraining tasks for modalities alignment.
6.1 Pretraining for modalities alignment
Many techniques have been proposed to align multiple modalities, leveraging the fact that the information is somehow redundant (when text and image is present), allowing to align the multiple modalities. The different cross-modal pretraining tasks in VrDU models aim to align textual, positional, and visual modalities, are summarized in Table 4.
Inspired by the masked language modeling (MLM) training approach in NLP, the Masked Visual Language Modeling (MVLM) was introduced by LayoutLM Xu et al. (2019), which masks text but not the layout information. Another possibility is to mask bounding boxes while keeping the text visible, as done in StructuralLM Li et al. (2021a). Joint text-layout (JTL) prediction Tang et al. (2023) and Area-Masked Language Modeling (AMLM) Hong et al. (2021) mask both text and layout. Masked Detection Modeling (MDM) Liao et al. (2023), Text-Layout Pretraining (TLP) Mao et al. (2024) and Layout Prediction (LP) Tang et al. (2023) predict masked text and layout using visual features.
Inspired by Masked Autoencoders (MAE) He et al. (2021a), which mask large portions of an input image and train the model to reconstruct the missing parts, several approaches have adapted this concept to capture finer visual and contextual details in document understanding. Masked Image Modeling (MIM) Huang et al. (2022) leverages discrete Variational Autoencoders (dVAE) to transform image patches into tokens, allowing token-level comparison to reconstruct masked regions. Building on this, Masked Image Reconstruction (MIR) Tang et al. (2023); Li et al. (2023e) directly predicts RGB pixel values for masked document regions. The Learn to Reconstruct (LTR) Appalaraju et al. (2021) task does the same for the whole image. Masked Region Modeling (MRM) Chen et al. (2020) incorporates both text and visible regions to infer the masked areas, while Image Fitting (IMG-FIT) Wu et al. (2021) shifts the focus to selecting the correct image from multiple options, providing a discriminative task.
Contrastive strategies train models to match or distinguish modality pairs by adjusting their similarity in a shared space. DoCo Li et al. (2024d) aligns text tokens and bounding boxes with visual features to connect words to their spatial context. LayoutLMv2 Xu et al. (2022) introduced Text-Image Alignment (TIA) to align text groups with image regions, using binary cross-entropy to predict if regions contain the text. LayoutLMv3 Huang et al. (2022) improved this with Word-Patch Alignment (WPA), aligning individual text tokens with image patches from a ViT. Region Alignment (WRA) Chen et al. (2020) uses Optimal Transport (OT) to align text tokens and image regions by finding a transport plan that matches word and visual region distributions, minimizing the alignment cost measured by cosine distance.
Captioning Strategies (IC) can be used to align text and image. Text Describes Image (TDI) task, used in DocFormer Appalaraju et al. (2021), aims at verifying if text matches a document image. Pramanik et al. (2022) later proposed Document Topic Modeling (DTM), using semantic contexts as soft labels to learn visual features and represent documents as topic distributions.
Finally, as layout does not translate to unambiguous reading order of the document, pretraining tasks linked to this aspect include Reading Order Prediction (ROP) Peng et al. (2022), training the model to recognize correct order between blocks, or Block-Ordering Predictions (B-ORD) Wu et al. (2021), which swaps two content blocks for the model to predict if the order is correct. Token-to-Grid (TTG) Appalaraju et al. (2023) divides the document into a grid and predicts which token belongs to which cell. Pramanik et al. (2022) extend reading-order tasks to multi-pages with Document Shuffle Prediction (DSP) task, training the model to detect shuffled pages in multi-page documents.
6.2 Pretraining Vision-only models
Hu et al. (2024a) highlight that vision-only models must be pre-trained on document parsing tasks to be capable of understanding the document content. To do so, it is possible to exploit the fact that documents are often generated from a symbolic source document (e.g. LaTeX, HTML), so the pretraining task aims at converting images into a structured text format: Webpage screenshots into HTML Lee et al. (2023a), PDF pages into LaTeX Wei et al. (2023) or Markdown Blecher et al. (2023), and table/chart into extended Markdown formats Hu et al. (2024a), CSV or JSON Kim et al. (2021). Hu et al. (2024a) implements a multi-format hierarchical reconstruction task named Unified Structured Learning, using multiple publicly available datasets and constructing structure-aware text sequences or text-position pairs as the targets.
We note that this pre-training task does not necessarily require an LLM decoder; to conserve resources, a lightweight decoder could be used during this phase Wei et al. (2024). However, most approaches leverage this pre-training task to align VrD features with the LLM decoder, training not only the visual encoder but also the projector Hu et al. (2024a).
Models Doc Info Chart DUDE MPDoc MMLong
VQA VQA QA VQA Bench
Multimodal (text, layout, visual) VrDU (Sec 2)
TILT 2021 87.1
LayoutLMv3 2022 83.4 45.1 20.3² 55.3²
ERNIE-Layout 2022 88.4
UDOP 2023 84.7 47.4
ViTLP 2024 65.9 28.7
DocFormerv2 2023 87.8 48.8 50.8² 76.8²
HiVT5 2023a 23.0 62.0
LayoutLLM 2024 86.9
DocLayLLM 2024 78.4 40.9
GRAM 2024 86.0 53.4 80.3
Arctic-TILT 2024 90.2 57.0 58.1 81.2 25.8
Vision-only VrDU models (Sec 3)
DONUT 2021 72.1 11.6 41.8
DESSURT 2022 63.2
Pix2Struct 2023a 76.6 40.0 58.6 62.0*
SeRum 2023b 77.9
LLaVAR 2023 6.73 12.3 13.5
Unidoc 2023 7.70 14.7 10.88
mPLUGDoc 2023b 62.2 38.2 57.4
QwenVL 2023a 65.1 35.4 65.7 84.4* 6.1*
UREADER 2023b 65.4 42.2 59.3
Monkey 2023 66.5 36.1 65.1 6.2*
TextSquare 2024 84.3 51.5 79.6
TextMonkey 2024 73.0 28.6 66.9
TextHawk 2024 76.4 50.6 66.6
DocPedia 2023a 47.8 15.2 46.9
CogAgent 2023b 81.6 44.5 68.4
Vary 2023a 76.3 66.1
mPLUGDoc1.5 2024a 82.2 50.7 70.2 6.9*
TinyChart 2024b 83.6
ILMXC24KHD 2024b 90.0 68.6 81.0 56.1* 76.9* 10.3*
Idefics2 2024 74.0 56.0* 7.0*
TokenPacker 2024c 70.0
mPLUGDoc2 2024d 80.7 46.4 70.0 46.8 69.4
Closed Models
GPT-4V 88.4 75.1 78.5
GPT-4o 92.8 85.5 54.0 67.0 42.8
Table 5: Average Normalized Levenshtein Similarity (ANLS) on single and multi-page VQA. ² denotes Single-page-native models concatenating page representations for multi-page; * denotes models using a retriever (PDF-Wukong Xie et al. (2024b) for InternLMXComposer2-4KHD, Naidu et al. (2024) for Pix2Struct, M3DocRAG Cho et al. (2024) for QwenVL and Idefics2). The top-3 scores are in bold.
7 Conclusion and Discussions
LLM-based approaches embedding the VrD representation directly into LLMs through mechanisms, like cross or composite attention improves efficiency, require less training Laurençon et al. (2024) while leveraging LLM capabilities for VrD features (see Section 5).
On the encoder side, we have two possibilities; encoding VrDs with its visual, textual and positional features, or in a vision-only approach. Although LLM-enhanced Vision-only methods (Section 3) are prominent in the recent literature, they face challenges in balancing coarse and fine-grained VrD representations without excessive computational complexity or compression, and are thus not even applicable to multi-page processing without a retriever (see Table 5). We think that approaches leveraging the three modalities (textual, visual and positional) of the document are more efficient, mainly on multi-page use-cases (see Table 5) but remain challenging due to the alignment requirement of those modalities. We believe that studying how models leverage these modalities to answer questions would be valuable to understand the role of each modality in generating responses, whether there are redundancies among them, if one modality dominates the others during generation, and whether they are truly complementary or not.
Based on the results in Table 5, models that maximize the use of positional information appear to achieve the best performance. This suggests that text and layout information alone might be sufficient to answer questions, even for complex charts and figures, provided the layout is optimally taken into account. Leveraging positional biases in attention—especially semantic-based relative positional biases like disentangled attention Peng et al. (2022) or Rich Attention Lee et al. (2022)—proves effective, particularly for tasks like Information Extraction Jaume et al. (2019); Park et al. (2019). Combining positional and semantic alignment of textual and visual modalities seems to allow great performances’ growth Borchmann et al. (2024).
We observed that there is still a challenge of multi-page documents. Retriever-based methods limit inter-page and inter-modality interactions, restricting their ability to perform cross-page reasoning, as we can see on MMLongBenchDoc Ma et al. (2024c) performance as depicted in Table 5. Emerging approaches that encode the document page-by-page with sparse-attention mechanisms for inter-page connections, like global-local or blockwise attention, represent a promising direction. These techniques have demonstrated superior performance in cross-page reasoning tasks Borchmann et al. (2024), positioning them as the future of multi-page document understanding. We also think that slicing approaches that use local information from cropped regions, combined with low-resolution global views Zhang et al. (2024d) could be leveraged for multi-page representation.
Assuming that documents will increasingly be born-digital, we believe the community should focus on models that process documents using text, layout, and optionally visual representation. These models are less costly, more scalable, and particularly more effective for handling multi-page documents.